### v1

@@ Abstract
We are concerned with the evolution of a density matrix $\rho(t)\in\mathbb{C}^{N\times N}$, which is a unit-trace symmetric positive semi-definite (SPSD) matrix, describing the state of a quantum system (e.g. a quantum computer), where $\rho$ is governed by the Lindblad, [1], equation
@@
To design CPTP schemes is a non-trivial task. For example, it is known, [3], that no explicit Runge-Kutta method is CP. Consequently, considerable effort has been spent on designing the numerical scheme that are on Kraus form. A pioneering work in this direction is [4], where a Taylor series expansion technique is used and explicit formulae defining schemes up to fourth order accuracy are provided. Later work [5, 6, 7] are based on various techniques including the use of Duhamel’s principle and principles of stochastic unraveling. While these aforementioned schemes are of interest we remark that they are single-step, single-stage methods, and the techniques used makes the construction of high order accurate schemes complicated.
Further, another property of the density matrix that is prominent in many quantum systems, in particular for the ones with low entropy, is that it can be effectively approximated by low rank techniques [8]. The low rank property is expected to be relevant for systems that are weakly coupled to the environment, or for the early stage of the dynamics of systems initialized in a pure state. When present, low rank structure can and should be exploited to accelerate simulations and reduce memory footprint. Moreover, as shown later in this paper, the Kraus form plays well with low rank structures and operations can be performed directly on low rank Cholesky factors. In the literature, a standard approach for evolving low rank structures is the time-dependent variational principle (TDVP) [9]. TDVP has been implemented in [10, 8] to approximate the time evolution of low rank density matrices. The main idea in TVDP is to evolve solution of an equation projected onto the low rank matrix manifold. However, since the tangent projection of the Lindbladian onto the low rank density matrix manifold in general may not be Lindbladian, a numerical scheme with the CP property resulting from TDVP is not possible to the best of the authors’ knowledge.
@@
Before defining the scheme, for reasons soon to be revealed, we consider the action of $e^{\mathcal{L}_{J}\Delta}$ onto a SPSD matrix $q(t)$. Such a matrix admits a Cholesky factorization $q(t)=V(t)V(t)^{\dagger}$. After one timestep, at time $t+\Delta t$, we have $q(t+\Delta t)=e^{\mathcal{L}_{J}\Delta t}q(t)$ which can also be represented as $q(t+\Delta t)=V(t+\Delta t)V(t+\Delta t)^{\dagger}.$ Here, as a consequence of the equivalence of
@@
We now present numerical examples illustrating the features of the new schemes. We denote by IF the scheme based on (9) with the flow operator $U(\cdot)$ approximated by matrix exponentiation. We denote by IF-LR and IF-LR-T the low rank implementation of the scheme based on (9) with the flow operator $U(\cdot)$ approximated by matrix exponentiation and the flow operator $U(\cdot)$ approximated by a Taylor series expansion, respectively. For all the results in this section we use the Butcher Tableau in (9) corresponding to the classic fourth order accurate Runge-Kutta method.
@@
In the left part of Figure 1 we display $\rho_{33}(t)$ computed using the IF method and the classic fourth order Runge-Kutta method, both using a timestep $\Delta t=0.1fs$. As can be seen in the inset the Runge-Kutta method does not stay CPTP while the IF method does. In the right part of Figure 1 we again display $\rho_{33}(t)$ obtained with the same methods but now with a five times larger timestep. For this larger timestep, not only is the loss of CPTP is more pronounced (especially at early times), but what is also clear is that the IF method is much more accurate for this example, despite both methods being of order four. In the right subfigure in Figure 1 we have also included results using IF-LR-T using a sixth order accurate Taylor series, yielding results that are very similar.
Figure 1: Evolution of the population $\rho_{33}$ in the example from [3]. To the left we compare the classic RK4 method (black) and the integrating factor method with matrix exponentiation for the flow (red, dashed) using a timestep of 0.1 femto-seconds. To the right we use a five times larger timestep and have also included results when the flow is computed using a Taylor series method of order 6.
@@
In summary, this paper develops a class of high order low rank CPTP schemes for the Lindblad equation with time-independent Hamiltonian and jump operators. We demonstrate in numerical examples that the schemes are accurate and efficient for problems with low rank properties. In subsequent work, we will consider the general case when the Hamiltonian and jump operators depend on time. We also plan to develop tensor network based low rank schemes [16, 17] to tackle the challenging case of many qubits.
Acknowledgment and Disclamer

### v2

@@ Abstract
We are concerned with the evolution of a density matrix $\rho(t)\in\mathbb{C}^{N\times N}$, which is a unit-trace symmetric positive semi-definite (SPSD) matrix, describing the state of a quantum system (e.g. a quantum computer), where $\rho$ is governed by the Lindblad equation [1]
@@
To design CPTP schemes is a non-trivial task. For example, it is known, [3], that no explicit Runge-Kutta method is CP. Consequently, considerable effort has been spent on designing the numerical scheme that are on Kraus form. A pioneering work in this direction is [4], where a Taylor series expansion technique is used and explicit formulae defining schemes up to fourth order accuracy are provided. Later work [5, 6, 7] are based on various techniques including the use of Duhamel’s principle and principles of stochastic unraveling. While these aforementioned schemes are of interest we remark that the techniques used when deriving them makes the construction of high order accurate schemes somewhat involved.
Further, another property of the density matrix that is prominent in many quantum systems, in particular for the ones with low entropy, is that it can be effectively approximated by low rank techniques. The low rank property is expected to be relevant for “systems that are weakly coupled to the environment, or for the early stage of the dynamics of systems initialized in a pure state” [8] (for further examples and motivation for low rankedness see [8]). When present, low rank structure can and should be exploited to accelerate simulations and reduce memory footprint. Moreover, as shown later in this paper, the Kraus form plays well with low rank structures and operations can be performed directly on low rank Cholesky factors. In the literature, a standard approach for evolving low rank structures is the time-dependent variational principle (TDVP) [9]. TDVP has been implemented in [10, 8] to approximate the time evolution of low rank density matrices. The main idea in TVDP is to evolve solution of an equation projected onto the low rank matrix manifold. However, since the tangent projection of the Lindbladian onto the low rank density matrix manifold in general may not be Lindbladian, a numerical scheme with the CP property resulting from TDVP is not possible to the best of the authors’ knowledge.
@@
Before defining the scheme, for reasons soon to be revealed, we consider the action of $e^{\mathcal{L}_{J}\Delta t}$ onto a SPSD matrix $q(t)$. Such a matrix admits a Cholesky factorization $q(t)=V(t)V(t)^{\dagger}$. After one timestep, at time $t+\Delta t$, we have $q(t+\Delta t)=e^{\mathcal{L}_{J}\Delta t}q(t)$ which can also be represented as $q(t+\Delta t)=V(t+\Delta t)V(t+\Delta t)^{\dagger}.$ Here, as a consequence of the equivalence of
@@
We now present numerical examples illustrating the features of the new schemes. We denote by IF the scheme based on (9) with the flow operator $U(\cdot)$ approximated by matrix exponentiation. We denote by IF-LR and IF-LR-T the low rank implementation of the scheme based on (9) with the flow operator $U(\cdot)$ approximated by matrix exponentiation and the flow operator $U(\cdot)$ approximated by a Taylor series expansion, respectively. For all the results in this section we use the Butcher Tableau in (9) corresponding to the classic fourth order accurate Runge-Kutta method. Note that we prefer to use non-dimensionalized quantities but that in order to repeat the numerical experiment in [3] we adopt their notation, which includes physical units.
@@
In the left part of Figure 1 we display $\rho_{33}(t)$ computed using the IF method and the classic fourth order Runge-Kutta method, both using a timestep $\Delta t=0.1fs$ (the same as in [3]). As can be seen in the inset the Runge-Kutta method does not stay CPTP while the IF method does. In the right part of Figure 1 we again display $\rho_{33}(t)$ obtained with the same methods but now with a five times larger timestep. For this larger timestep, not only is the loss of CPTP is more pronounced (especially at early times), but what is also clear is that the IF method is much more accurate for this example, despite both methods being of order four. In the right subfigure in Figure 1 we have also included results using IF-LR-T using a sixth order accurate Taylor series, yielding results that are very similar.
Figure 1: Evolution of the population $\rho_{33}$ in the example from [3]. The units for the $x$-axis is pico seconds. The number of timesteps in the left and right insets are 300 and 60 respectively. To the left we compare the classic RK4 method (black) and the integrating factor method with matrix exponentiation for the flow (red, dashed) using a timestep of 0.1 femto-seconds. To the right we use a five times larger timestep and have also included results when the flow is computed using a Taylor series method of order 6.
@@
In summary, this paper develops a class of high order low rank CPTP schemes for the Lindblad equation with time-independent Hamiltonian and jump operators. We demonstrate in numerical examples that the schemes are accurate and efficient for problems with low rank properties. In subsequent work, we will consider the general case when the Hamiltonian and jump operators depend on time. We also plan to develop tensor network based low rank schemes [16, 17] to tackle the challenging case of many qubits. As pointed out by one of the reviewers, the low rank techniques used here can likely be adopted to other methods, for example those in [4, 5].
Acknowledgment and Disclaimer
@@
This material is based upon work supported by the National Science Foundation under Grant Number 2436319. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.